OpenStudy (anonymous):

Ashok starts to walk from the point P. He went 10 km towards North and 6 km toward South. After this he went 3 km towards East. How far is he now from P?

OpenStudy (anonymous):

This is really easy, where are you stuck?

OpenStudy (anonymous):

Anyways, (3,4,5) is a Pythagorean triple.

Parth (parthkohli):

 Notice how it forms a right-angled triangle. We can find the value by using Pythagorean theorem. Notice that the required side is the Hypotenuse. \(\Large \color{purple}{\rightarrow ?^2 = 3^2 + 4^2 }\) \(\Large \color{purple}{\rightarrow ?^2 = 9 + 16 }\) \(\Large \color{purple}{\rightarrow ?^2 = 25 }\) \(\Large \color{purple}{\rightarrow ? = \sqrt{25} }\) \(\Large \color{purple}{\rightarrow ? = 5 }\) \( \checkmark \)